OpenStudy (anonymous):

solve the system by using elimination or elimination with multiplication.One point for steps and solution. x-y=3 and x+y=7

OpenStudy (anonymous):

Subtract the first equation from the second. \[x+y=7\]-\[x-y=3\]\[=2y=4\]Divide by 2 \[y=2\]Plug back 2 in for y into either equation. \[x-2=3\]\[x=5\] x=5, y=2

OpenStudy (anonymous):

ty